### v1

# An Engorgio Prompt Makes Large Language Model Babble on

Abstract
Auto-regressive large language models (LLMs) have yielded impressive performance in many real-world tasks. However, the new paradigm of these LLMs also exposes novel threats. In this paper, we explore their vulnerability to inference cost attacks, where a malicious user crafts Engorgio prompts to intentionally increase the computation cost and latency of the inference process. We design Engorgio, a novel methodology, to efficiently generate adversarial Engorgio prompts to affect the target LLM’s service availability. Engorgio has the following two technical contributions. (1) We employ a parameterized distribution to track LLMs’ prediction trajectory. (2) Targeting the auto-regressive nature of LLMs’ inference process, we propose novel loss functions to stably suppress the appearance of the <EOS> token, whose occurrence will interrupt the LLM’s generation process. We conduct extensive experiments on 13 open-sourced LLMs with parameters ranging from 125M to 30B. The results show that Engorgio prompts can successfully induce LLMs to generate abnormally long outputs (i.e., roughly 2-13$\times$ longer to reach 90%+ of the output length limit) in a white-box scenario and our real-world experiment demonstrates Engergio’s threat to LLM service with limited computing resources. The code is accessible at https://github.com/jianshuod/Engorgio-prompt.
1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Ouyang et al., 2022; Carlini et al., 2021) have demonstrated remarkable performance in various real-world applications, e.g., online chatting (Shen et al., 2023), customer service (Gimpel et al., 2023), and finance (Wu et al., 2023). Given the increasing popularity and adoption of LLMs, reducing their inference cost becomes critical. Firstly, from the cost aspect, a modern LLM normally contains billions of parameters, and each inference generation may consume considerable resources and time. Many AI service providers are paying more bills to support their LLM inference services than training (Patel et al., 2024; Li et al., 2024; Patterson et al., 2022). Secondly, from the service availability aspect, there is fierce competition across different LLM service providers, making service reliability and fast response time important factors in attracting customers. Meanwhile, these two considerations motivate malicious entities to attack the LLMs, increasing their operational cost and generation latency.
In this paper, we explore the landscape of inference cost attacks against modern LLMs. First proposed in Shumailov et al. (2021) to attack encoder-decoder transformers, inference cost attacks aim to intentionally maximize the energy consumption and latency of model inference via a new type of adversarial input. The inference cost attacks on language models (Shumailov et al., 2021; Chen et al., 2022; Feng et al., 2024) are tailored for encoder-decoder models and rely on perturbation-based mutation to progressively hit a desirable adversarial input. However, as demonstrated in Section 2.2 and Section 4, they become ineffective against modern LLMs (Graves, 2013), which adopt the auto-regressive generation scheme (Graves, 2013), remove the cross-attention mechanism, and employ a sub-word tokenization algorithm. Geiping et al. (2024) propose an adversarial prompt attack to coerce LLMs into repeating specific content, achieving effects similar to an inference cost attack. However, its reliance on the starting response weakens robustness and increases detectability.
In general, it is challenging to design inference cost attacks against modern decoder-only LLMs, even given the existing works discussed above. The main challenges stem from two aspects: (1) Uncertain Generation Process. The generation process of decoder-only LLMs is inherently casual, auto-regressive, and sampling-based, rendering it difficult to constrain them to generate a specific long response. The occurrence of one deviant token can directly distort the generation process from the desirable decoding trajectory, challenging attack effectiveness and stability. (2) Discrete Input Modality. Text-completion LLMs accept input text in the form of discrete token sequences but operate within the embedding space, which implies an irreversible mapping from the embedding space back to the token space. While we can leverage gradient information to optimize a more desirable soft embedding representation for the input, we face challenges in accurately identifying corresponding tokens in the token space for the optimized soft embeddings. This restricts us from effectively leveraging gradients to guide updates to the input token sequence (i.e., adversarial input). To address the above challenges, we need to consider two intriguing and important questions: (1) how to accurately frame our goal as a well-aligned optimization problem and (2) how to effectively instruct the updates to the discrete input sequence given the modeled objective.
Figure 1: Distributions of the total lengths (input plus output) of normal samples from ShareGPT (sharegpt, 1) and Engorgio prompts.
In this paper, we introduce Engorgio [^1] Engorgio is a spell in the Harry Potter universe, which causes objects (or creatures) to increase in size., a simple yet effective method to generate threatening Engorgio prompts against state-of-the-art LLMs. Our focus is on text completion, where LLMs predict the next token based on the initial prompt and previously generated tokens until an end-of-sequence (<EOS>) token is predicted or a maximum length is reached. Technically, Engorgio effectively addresses the above challenges via: (1) Inspired by the special role of <EOS> token in determining whether the model halts its response, we adopt an untargeted objective called <EOS> escape loss, which reduces the <EOS> token’s occurrence probability. We also combine a self-mentor loss to stably induce longer responses. (2) We employ a re-parameterization design to effectively utilize the gradients, by modeling the potential distribution of the entire context that can fulfill both objectives. Figure 1 shows the effects of our attack: normal prompts (e.g., the renowned ShareGPT dataset (sharegpt, 1)) typically tempt the LLMs to produce short sequences; in contrast, the crafted Engorgio prompts can make the model response extraordinarily long.
In summary, our main contributions lie in three folds: 1) We explore a novel research direction, inference cost attack against modern auto-regressive LLMs. We highlight how crafted adversarial prompts can impact LLM service availability. 2) We analyze technical challenges associated with the attack surface. Based on our insights, we propose Engorgio, a simple yet effective method that can stably induce lengthy LLM responses. 3) To prove the effectiveness of Engorgio, we conduct extensive experiments over 6 base models and 7 supervised fine-tuned (SFT) models with parameters ranging from 125M to 30B, as listed in Table 5. Specifically, the generated Engorgio prompts can achieve roughly 90%+ of the maximum allowable length on 6 base models, while normal queries can only cause between 0-40%. For SFT models, Engorgio can significantly outperform baselines by up to 13$\times$. A real-world experiment demonstrates Engorgio’s implications in service availability.
2 Preliminaries
2.1 Large Language Models (LLMs)
The task of language modeling tracks the rationality of text sequences and treats the probability of a certain sequence as a product of conditional probabilities (Jelinek, 1980; Bengio et al., 2003):
$P(x_{1},\cdots,x_{N})=\prod_{i=1}^{N}P(x_{i}|x_{1},\cdots,x_{i-1}),$ (1)
where $P(x_{i}|x_{1},\cdots,x_{i-1})$ denotes the probability of predicting $x_{i}$ as the next token given the sequence $x_{1}\cdots x_{i-1}$. For a Transformer-based model $f_{\Theta}:\mathcal{X}\to\mathcal{Y}$, it accepts a sequence of tokens with any admitted length $S$ and produces an output vector $r_{S}=f_{\Theta}(X_{1:S})\in\mathbb{R}^{V}$ to predict the next token, where $V$ is the model’s vocabulary size. Most prevalent LLMs like LLaMA (Touvron et al., 2023a) and GPT-4 (OpenAI, 2023) are based on the Transformer decoder architecture (Vaswani et al., 2017). The architecture is designed to perform inference in an auto-regressive manner (Graves, 2013), i.e., LLMs generate one token at a time and use the previously generated tokens to instruct the prediction of the next token. We detail the LLM generation process and models involved in this work in Appendix A.1.
2.2 Inference Cost Attacks
Machine learning services are facing an availability threat. Shumailov et al. (2021) showed that malicious users could intentionally craft adversarial inputs, known as sponge examples, to significantly increase the energy consumption and latency of the corresponding inference process. Such inference cost attacks can greatly affect the service provider’s operational cost and user experience. Following this work, a variety of attacks have been designed to target different AI systems and applications, for example, image classification (Müller & Quiring, 2024), camera-based object detection (Shapira et al., 2023; Schoof et al., 2024; Shapira et al., 2022; Xiao et al., 2024; Ma et al., 2024), LiDAR-based object detection (Liu et al., 2023), and multimodal models (Gao et al., 2024).
This paper focuses on attacking the modern auto-regressive LLMs. Existing inference cost attacks against language models (Shumailov et al., 2021; Chen et al., 2022; Feng et al., 2024) are only effective when targeting the encoder-decoder structure. Shumailov et al. (2021) generated sponge examples by compressing more tokens into one sentence, leading to a higher computational burden in the cross-attention operations, which are ineffective for LLMs lacking cross-attention modules. Sub-word tokenization methods BPE (Sennrich et al., 2016) eliminate the appearance of <UNK> token and enhance LLMs’ typo-tolerating ability, largely invalidating perturbation-based methods like LLMEffiChecker (Feng et al., 2024). For the optimization-based method, Geiping et al. (2024) proposes a targeted attack that coerces LLMs into producing elicit a specific starting response (i.e., repeating ”Hello There” 24 times), indirectly achieving effects similar to an inference cost attack. However, this approach is less stable due to its reliance on the starting response’s effectiveness and is easily detectable as the starting response serves as a clear indicator of adversarial intent. This motivates us to design a new attack methodology tailored for modern LLMs. In this work, we propose a simple yet effective method to overcome the technical challenges inherent in this task.
2.3 Threat Model
We design our attack following the threat model of previous inference cost attack studies against language models (Shumailov et al., 2021; Chen et al., 2022; Feng et al., 2024) and provide detailed discussion about the practicality and implications of the attack in Appendix A.2.
  • $\bullet$
    Attacker’s goal: As a service user, the attacker aims to craft Engorgio prompts $\mathcal{T}$, which could induce as long output as possible. Such behaviors could bring much higher operational costs for the LLM service provider, and affect the service availability to other normal users.
  • $\bullet$
    Attacker’s knowledge: We mainly consider a white-box scenario, where the attacker has full knowledge of the target model, including its architecture, input template, model parameters, etc. We also consider the black-box setting, in which we transfer Engorgio prompts to attacker-unknown models (see Appendix B.1 for details).
  • $\bullet$
    Attacker’s capability: The attacker locally generates the Engorgio prompts $\mathcal{T}$, aligned with her knowledge settings. Then she sends the constructed Engorgio prompts $\mathcal{T}$ to the target LLMs and collects the responses for attack checking.
3 Methodology
3.1 Attack Insight and Overview
In order to achieve the attack goal, we review the mechanism of generating texts by LLMs. A sample for the LLM can be split into an input part and an output part (see Appendix A.3 for more analysis). Given an input sequence (dubbed prompt) composed of $k$ tokens, the model generates the subsequent tokens (i.e., the output part). The generation continues until either of two conditions is met: (1) reaching a pre-set maximum allowable length; (2) encountering an <EOS> token which indicates the end of the sentence. As the maximum allowable length is fixed as $S$, the problem is stated as follows: the later an <EOS> token is encountered in the inference process, the higher cost and latency this query will take. Therefore, to achieve latency damages to the service provider, i.e., maximizing the length of the output part, the attacker aims to create Engorgio prompts, which can effectively suppress the possibility of predicting the <EOS> token during the inference.
Based on this insight, we design Engorgio, a novel attack framework to generate Engorgio prompts against LLMs. Figure 2 shows its overall pipeline and we provide a term list in Appendix A.4. The core is the introduction of a parameterized proxy distribution. To satisfy the requirements for Engorgio prompts, we explore how to update the distribution with the guidance of an <EOS> escape loss and self-mentor loss. The whole process of crafting Engorgio prompts is two-stage:
  • $\bullet$
    Generation stage: For each optimization step, we convert the proxy distribution matrix $\theta$ to a weight matrix $w$ using the Gumbel-Softmax function. We then aggregate the embeddings of all token candidates weighted by $w$ to project $\theta$ into the embedding space. This output is fed into the model to calculate two loss terms, allowing us to obtain gradients for $\theta$ easily. The matrix $\theta$ is updated based on these losses, continuing until no significant changes are detected.
    Figure 2: The pipeline of Engorgio. The whole pipeline is divided into two stages. During the generation stage, we employ a gradient-based method to update the proxy distribution for the Engorgio prompt, where the gradient information is obtained from a local proxy model. For the testing stage, we leverage the optimized proxy distribution to decide the final Engorgio prompt, which is expected to escape the occurrence of the <EOS> token.
  • $\bullet$
    Testing stage: The optimization process guarantees that the output part falls onto a region with low probabilities of <EOS>. Given the strong correlation between the Engorgio prompt and the output, we can sample the Engorgio prompt using the normalized $w_{1:t}$. It is observed that as the optimization progresses, the distribution matrix $\theta$ typically converges toward a specific prompt with a significantly higher sampling probability compared to others. This prompt is adopted as the final Engorgio prompt $\mathcal{T}$. This approach significantly reduces the cost of evaluating other prompt candidates. We hypothesize that the objectives, particularly self-mentor loss, contribute to identifying the optimal Engorgio prompt, as detailed in Section 4.6.
3.2 Proxy Distribution
To increase the lengths of the target LLM’s responses, we search for the Engorgio prompts $\mathcal{T}$ with the help of a proxy model. LLMs typically accept a token sequence (corresponding to one input text) as input, cast each token into the embedding space, and work within the embedding space. Each token has a corresponding embedding; however, not all embeddings correspond to tokens. We can optimize suitable embedding expressions that satisfy the objectives in the form of prompt learning (Li & Liang, 2021; Liu et al., 2021), but we face challenges in determining the corresponding token sequence (i.e., input text). We resort to a re-parameterization method.
As LLMs predict the next tokens according to the probability distribution, it is more efficient to search for desirable Engorgio prompts by sampling from an appropriate distribution (Guo et al., 2021). Therefore, we introduce a proxy distribution to track the process of sequence sampling. This proxy distribution is parameterized as a matrix $\theta\in\mathbb{R}^{S\times V}$, with $S$ denoting the maximum allowable length, corresponding to the whole context. It instructs how to select a suitable token sequence from a token vocabulary with $V$ token candidates in the following test stage. Then the question is how to ensure that the proxy distribution $\theta$ meets the objectives. This endeavor seeks to involve the distribution matrix in the generation stage, subsequently updating it based on the gradients. Concretely, in the forward pass, the distribution vector $\theta_{i}$, corresponding to the $i$-th token in the Engorgio prompt where $i\in\left\{1,\cdots,S\right\}$, is independently normalized. This serves as a weight to aggregate token embeddings across the model vocabulary, thereby casting $\theta_{i}$ as a soft token in the embedding space. This process is formulated as Eq. 2.
$\tilde{e}(\theta_{i})=\sum_{j=1}^{V}(w_{i})_{j}e(j),$ (2)
where $e(j)\in\mathbb{R}^{H}$ denotes the embedding of the $j$-th token within the model vocabulary and $w_{i}\in\mathbb{R}^{V}$ is the normalized version of $\theta_{i}$ with the sum $\sum_{j=1}^{V}(w_{i})_{j}=1$. We adopt Gumbel-Softmax (Jang et al., 2017), which introduces stochastic elements and enriches the diversity of tokens involved in the generation stage. The normalization is conducted in Eq. 3:
$(w_{i})_{j}=\tfrac{\text{exp}((\theta_{i,j}+g_{i,j})/\tau)}{\sum_{k=1}^{V}\text{exp}((\theta_{i,k}+g_{i,k})/\tau)},$ (3)
where $g_{i,1}\cdots g_{i,V}$ are drawn from the distribution Gumbel(0,1) and $\tau>0$ is a temperature factor used to control the uncertainty. The introduction of the random variable $g_{i,k}$ from an i.i.d distribution benefits the diversity of the sampling operation. Due to the differentiability of Gumebl-Softmax, we can take full advantage of the gradient information to update $\theta$ in the generation stage and guide the sampling of the final Engorgio prompt $\mathcal{T}$ in the test stage.
SFT models assume the input should be embedded in a specified template $T$, as illustrated in Figure 6. Considering the most general case that the template $T=\left\{\left[P_{1:i}\right],x,\left[P_{i+1:m}\right],y\right\}$ contains a prefix and an infix, we define the corresponding embedding sequence to $\theta$ as Eq. 4.
$E(\theta)=\left\{e(\left[P_{1:i}\right]),\tilde{e}(\theta_{1:t}),e(\left[P_{i+1:m}\right]),\tilde{e}(\theta_{t+1:s-m})\right\},$ (4)
where $P_{1:i}$ and $P_{i+1,m}$ represent the token sequences corresponding to prefix and infix, and the shape of $\theta$ is adjusted to $(s-m)\times V$. The input composition is illustrated in Figure 6.
3.3 Loss Design
To obtain a desirable proxy distribution, we mainly depend on two key loss components to update $\theta$, <EOS> escape loss and self-mentor loss. The <EOS> escape loss closely aligns with our target goal to make the output part longer while the self-mentor loss is designed to enhance the usability of the proxy distribution. Balancing the impact of the two loss terms with $\lambda$, we update the proxy distribution as follows:
$\min_{\theta}\ \mathcal{L}_{esc}(\theta)+\lambda\mathcal{L}_{sm}(\theta)$ (5)
<EOS> escape loss. Due to the unpredictability of the LLM generation process, enforcing a specified long response is challenging. We resort to an untargeted objective, which is to decrease the prediction probability of the <EOS> token. However, it is still impossible to accurately forecast the exact occurrence position of <EOS> during the test stage. To tackle this, we propose penalizing the occurrence of <EOS> token from all positions, rather than focusing on specific positions. This broader treatment allows us to effectively manage the uncertainties associated with <EOS> placement. The <EOS> escape loss is defined as below:
$\mathcal{L}_{esc}(\theta)=\sum_{i=1}^{S}\text{Softmax}(f_{\Theta}(E(\theta)_{1:i}))_{\kappa},$ (6)
where $\kappa$ denotes the index of the <EOS> token for the target model. We adopt a Softmax-normalized probability of <EOS> so that it can better measure the relative chance that the model predicts <EOS> as the next token at a certain position, which is more effective than directly decreasing the absolute logit of the <EOS> token. An input sequence containing <EOS> is illegal, as the inference process should have halted before predicting the next tokens for the Engorgio prompt. Therefore, we also consider reducing the predicted <EOS> probabilities of the Engorgio prompt part.
Self-mentor loss. Another challenge is that we can only query the target model utilizing the Engorgio prompt $\mathcal{T}$ to ensure attack stealthiness and efficiency. Considering the auto-regression nature of modern LLMs, we cut off the first $t$ tokens as our Engorgio prompt. Moreover, $\theta_{i}$ only independently tracks the token selection of the $i$-th position, but the correlation between tokens should also be enhanced. Therefore, we seek to enhance the relevance of all tokens in the sequence, especially the bond between the Engorgio prompt and output parts. Inspired by LLM’s causal pre-training paradigm, we search for a sequence where the proxy model fits well. The loss term is given below:
$\mathcal{L}_{sm}(\theta)=\sum_{i=1}^{S}\mathcal{L}(w_{i+1},\text{Softmax}(f_{\Theta}(E(\theta)_{1:i}))),$ (7)
where $\mathcal{L}$ is the cross entropy loss. The closer to 0 $\mathcal{L}_{sm}$ is, the better the proxy model fits in input $E(\theta)_{1:S}$, which helps the Engorgio prompt $\mathcal{T}$ steadily induce a longer output.
4 Evaluation
4.1 Experimental Setup
LLMs. We include multiple base models, OPT-125M (Zhang et al., 2022), OPT-1.3B (Zhang et al., 2022), GPT2-large (Radford et al., 2019), LLaMA-7B (Touvron et al., 2023a), LLaMA-2-7B (Touvron et al., 2023b), and LLaMA-30B (Touvron et al., 2023a). SFT models are further fine-tuned with additional dataset, for which we consider seven well-known SFT models including Alpaca (7B) (Taori et al., 2023), Vicuna (7B) (Chiang et al., 2023), StableLM (7B) (stablelm, 2), Koala (7B) (koala, 3), Oraca (7B) (Mukherjee et al., 2023), Samantha (7B) (samantha, 4), and ChatGLM (6B) (Du et al., 2022), as our targets. Considering the crucial importance of prompts, we also consider the three cases of deploying base models with prompts according to the attacker’s knowledge about the deployed prompt. Constrained by space limit, the corresponding results are presented in Appendix B.3.
Baselines. We consider four types of inputs as baselines for comparisons. (1) Normal inputs: we collect 50 samples from the training dataset for Standford-alpaca (Taori et al., 2023), which are generated by OpenAI’s text-davinci-003, and 50 samples from ShareGPT (sharegpt, 1), a website where people can share their ChatGPT conversations. We use the mixup to roughly represent the normal response length of LLMs. (2) Special inputs: we use prompts with the semantics of demanding a longer output ($i.e.$, prompts starting with “output longer”). (3) LLMEffiChecker: we adopt the three attacks (character, word, and structure attack) proposed in Feng et al. (2024) and report the averaged results across the attack variants. (4) Sponge examples: we generate such inputs using the method from Shumailov et al. (2021) by only setting the same input length as our method.
Metrics. Due to the intractable serving mechanisms for LLM, we report results on the level of model behaviors. To mitigate potential sampling bias caused by the inherent variability in LLM inference, we measure the average token number of the generated outputs (Avg-len). We query the target LLM multiple times using the sampling generation and compute the average length across these responses. This renders Avg-len a robust estimate of the Engorgio prompt’s efficacy. Second, we calculate the ratio of the LLM outputs that reach the maximum length (Avg-rate) to evaluate the stability. Notably, inference costs increase super-linearly with longer responses, making Avg-len a lower bound on the prompt’s impact on inference cost. Further analysis of the Engorgio prompt’s impact is detailed in Appendix A.5.
Configurations. We use the Adam optimizer with a learning rate of 0.1 to update the distribution matrix $\theta$. We allow a maximum of 300 optimization steps, the cost of which is acceptable, especially when considering the reusability as explained in Appendix A.6. The Gumbel-Softmax temperature factor $\tau$ is set to 1, and the default Engorgio prompt length is $t=32$. The input length of normal inputs, special inputs, LLMEffiChecker, and sponge examples is roughly the same as Engorgio to ensure fairness. According to our extensive experiments, the loss coefficient $\lambda$ is empirically set to 1. The optimization starts with a random prompt and we use it to initialize the proxy distribution. Constrained by the computing resources, we set 1,024 as the pre-set maximum length. We also conduct experiments with full context size on two representative base models (2,048 for LLaMA-30B and LLaMA-7B) and one SFT model (4,096 for Samantha) to demonstrate the extensibility to longer context size. Please refer to Appendix B.8 for examples of prompts and responses for normal inputs, sponge examples, and Engorgio prompts.
4.2 Main Results
We report our results on base models in Table 1. Comparing normal and special inputs reveals that semantic inputs induce base models to output longer. This means that base LLMs can understand the semantics inside the inputs and seemingly feature being talkative. However, relying solely on special input is far from reaching the maximum allowable length. LLMEffiChecker proves ineffective against more advanced LLMs. Our method can achieve a very high ratio (roughly 90-100%) of letting the base model keep endlessly generating tokens until reaching its maximum length, which outperforms all baselines including sponge examples. While sponge examples extend output length compared to normal or special inputs, they are less stable than Engorgio as they struggle with LLMs’ sampling-based decoding. Results of more base models are presented in Appendix B.2.
Table 1: Results of Engorgio against modern LLMs.
  Model LLaMA-30B LLaMA-7B LLaMA-7B LLaMA-2-7B
Max length 2048 1024 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 622.2 12% 611.4 39% 757.9 16% 818.9 69%
Special inputs 1005.2 40% 737.3 58% 1292.2 54% 773.8 50%
LLMEffiChecker 1052.8 28% 682.3 41% 1306.7 41% 833.7 64%
Sponge examples 1277.7 38% 857.6 81% 1659.8 78% 900.6 86%
Engorgio 2019.1 95% 983.4 94% 1817.7 84% 1024 100%
  Model Samantha StableLM Koala Orca
Max length 4096 1024 1024 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 313.7 2% 388.2 6% 357.5 6% 286.0 1%
Special inputs 173.3 0% 202.9 4% 436.1 0% 199.5 0%
LLMEffiChecker 172.6 0% 688.3 38% 324.6 5% 203.4 0%
Sponge examples 284.9 3% 301.1 16% 432.2 22% 211.1 1%
Engorgio 3951.5 95% 1021.6 98% 1024 100% 908.1 86%
Prefix+Engorgio 4027.6 95% 1024 100% 1024 100% 962.6 90%
  Model Samantha ChatGLM Alpaca Vicuna
Max length 1024 1024 1024 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 231.6 2% 263.4 1% 179.4 0% 312.6 0%
Special inputs 82.0 0% 247.9 0% 132.5 7% 252.4 4%
LLMEffiChecker 149.8 0% 182.0 0% 192.9 5% 273.4 0%
Sponge examples 155.1 0% 685.2 56% 833.8 78% 599.6 44%
Engorgio 944.0 89% 979.6 95% 954.2 92% 789.3 60%
Prefix+Engorgio 970.0 89% 1024 100% 1024 100% 861.5 68%
SFT models may use cut-off as a preprocessing strategy on their fine-tuning datasets (e.g., at most 512 tokens for Alpaca). This potentially biases the fine-tuned model to produce short responses, which makes our goal challenging, as suggested by the results of normal inputs in Table 1. For special inputs, even with instructions for longer responses, SFT models still produce notably shorter outputs, sometimes even shorter than normal inputs. The silent nature of SFT models worsens the performance of sponge examples. For LLMEffiChecker, the weaker performance extends to SFT models. We hypothesize that recent LLMs are more robust to typing errors, invalidating perturbation-based attacks. In contrast, Engorgio knows how to better optimize the Engorgio prompt by focusing on a distinct goal: avoiding the generation of the <EOS> token. It effectively increases the output length to approach the maximum limit, especially when paired with a semantic prefix as discussed in Section 4.4, achieving near-maximum allowable lengths. We also explore a black-box scenario, where we resort to the transferability of Engorgio prompts (See Appendix B.1) for details.
4.3 Ablation Study
Impact of loss design. We explore the efficacy of the two loss terms. Initially, we assess the performance when optimizing only with the <EOS> escape loss (noted as “ESC” in Table 3). A comparison with normal input and special input from Table 1 reveals that even utilizing only the <EOS> escape loss consistently results in longer outputs. We also observe that combining the self-mentor loss, identified as “ESC+Self-mentor” in Table 3, further increases the Avg-len with almost no extra cost. More experiments in Appendix B.5 show that the performance of Engorgio is not strictly dependent on the choice of $\lambda$.
Impact of Engorgio prompt length. We explore the attack results under different prompt lengths $t$. The basic intuition is that a longer Engorgio prompt can contain more malicious information to induce the LLMs’ outputs to be longer. The results are given in Table 3 with three different prompt lengths (i.e., 32, 64, and 128). For base models like OPT-125M and LLaMA-7B, even the smallest prompt length of 32 can induce them to output max-length sequences. For SFT models, as the prompt length increases, Avg-len and Avg-rate increase in most cases. In summary, although a longer prompt improves the attack performance, it is not a prerequisite for ensuring effectiveness.
Table 2: Ablation study. Prompt length is separated from Avg-len to better understand the impact of key designs.
  LLaMA-7B (1024)
Prompt length ESC ESC+Self-mentor
Avg-len Avg-rate Avg-len Avg-rate
  32 893.5 + 32 80% 951.4 + 32 94%
64 870.6 + 64 85% 945.5 + 64 98%
128 827.8 + 128 85% 880.6 + 128 94%
  Alpaca (1024)
Prompt length ESC ESC+Self-mentor
Avg-len Avg-rate Avg-len Avg-rate
  32 967.4 + 32 96% 992.0 + 32 100%
64 943.8 + 64 96% 949.1 + 64 98%
128 867.3 + 128 96% 896.0 + 128 100%
  Koala (1024)
Prompt length ESC ESC+Self-mentor
Avg-len Avg-rate Avg-len Avg-rate
  32 980.0 + 32 98% 992.0 + 32 100%
64 950.5 + 64 98% 960.0 + 64 100%
128 849.2 + 128 90% 896.0 + 128 100%
Table 3: Impact of temperature setting.
                           StableLM (1024)
Temperature Avg-len Avg-rate
  0.1 1021.6 98%
0.3 830.5 62%
0.5 610.4 28%
0.7 513.8 33%
                           Samantha (1024)
Temperature Avg-len Avg-rate
  0.1 944.1 90%
0.3 714.1 58.7%
0.5 553.0 40%
0.7 406.3 23.8%
                           ChatGLM (1024)
Temperature Avg-len Avg-rate
  0.1 979.6 95%
0.3 934.0 88%
0.5 908.1 81%
0.7 820.0 71%
4.4 Attacks at Different Decoding Temperatures
We investigate how temperature affects Engorgio’s effectiveness. Results in Table 3 show that a larger temperature introduces more uncertainty during generation, potentially leading to deviations in the model response. For talkative base models, they are tempted to respond endlessly when a high temperature of 0.7 is used while a low temperature of 0.1 is more suitable for the silent SFT models. In most cases, e.g., when querying API service, the temperature is at the users’ discretion. Our quantitative statistics show that the output lengths induced by Engorgio prompts gather either at the shorter end or around the maximal length. See details in Appendix B.4. Engorgio prompts can either encourage the SFT model to generate longer outputs or confuse it, resulting in brief responses like ”not understand.” Thus, we consider fusing Engorgio prompts with semantic instructions.
Table 4: Results of introducing semantic prefix/suffix when the temperature is 0.7.
  Alpaca StableLM
Model 1024 1024
Avg-len Avg-rate Avg-len Avg-rate
Only prefix 132.5 7% 202.9 4%
Prefix+normal 214.2 1% 440.6 9%
Engorgio 353.2 8% 513.8 33%
Prefix + Engorgio 531.2 43% 884.8 83%
Engorgio + suffix 314.9 18% 534.3 38%
Adding semantic prefix/suffix. We can introduce additional semantic instructions to avoid the SFT model directly outputting “not understood”. Particularly, a prefix with semantics that encourages longer response will be woven with the Engorgio prompt in both the generation and testing stage. The results in Table 4 show that introducing semantic prefixes can improve performance. Comparing with the results in Table 1, we can observe that adding the prefix to normal inputs still cannot induce an extremely long response. Compared to adding a prefix, adding the same semantic sequence as a suffix does not help. We hypothesize that that’s because a semantic prefix impacts the entire generation process while a suffix only influences the subsequent generation.
4.5 Attacking Real-world LLM Services
We conduct a real-world case study to assess the practical threats of Engorgio. Corresponding to realistic scenarios listed in Appendix A.2, users share limited cloud resources for inference requests.
Experiment setup. We utilize the Hugging Face inference endpoint [^2] https://ui.endpoints.Huggingface.co/ as our cloud service, deploying StableLM (maximal length of 4096) as the target LLM. Our experiments explore three GPU configurations: 1× Nvidia A10, 4× Nvidia A10, and 2× Nvidia A100, aiming to demonstrate how a small number of attackers can significantly compromise the service’s performance. We focus on two main metrics: normal client latency, defined as the average response time from querying the service to receiving the output, and server throughput, calculated as the number of requests processed per minute. More details can be referred to in Appendix B.6.
Figure 3: Results of attacking real-world LLM services (“MU”: malicious user, “NU”: normal user).
Main results. As shown in Figure 3, attackers with Engorgio prompts can significantly compromise the LLM services. Although the inference time for normal clients remains consistent, Engorgio prompts significantly increase the queuing time for normal clients scheduled after the attackers. We observe that only a small ratio of attackers (e.g., 1 out of 10 or 5 out of 100) could lead to a significant latency increase. Besides the negative effect on clients, Engorgio also severely harms the cloud service throughput, which is almost cut off. We conclude that a limited number of attackers equipped with Engorgio could severely disturb the fragile cloud-based LLM services.
4.6 Why is Our Method Effective?
Q1. How does the distribution matrix instruct the token selection in the optimization process?
In the optimization process, we seek to update the distribution matrix $\theta$ rather than selecting individual tokens, meaning that the most suitable input is the aggregated embedding $\tilde{e}(\theta_{i})$. In Figure 5, we find that even in the middle of the optimization process, the token sequence greedily sampled according to the distribution matrix performs only slightly worse than the aggregated embeddings when our goal is to decrease the total loss. This means that our distribution update design is effective in searching for suitable token sequences. Moreover, we find that the <EOS> escape loss of SFT models like Koala is much harder to decrease than base models like OPT-125M. This partially supports that base models are easier to induce than SFT models.
Figure 4: Loss curves on OPT-125M (base model) and Koala (SFT model), with aggregated embeddings and token sequence as input, respectively.
Figure 5: On LLaMA-7B, the <EOS> escape loss correlates with the relative level of <EOS> being predicted.
Q2. Does <EOS> escape loss stop <EOS> from appearing?
To verify that <EOS> escape loss reduces the probability of <EOS> appearance, we calculate the highest probability of the <EOS> token at all the $S$ positions. We formulate this as $\mu=\max(\{{\text{Softmax}(f_{\Theta}(E(\theta)_{1:i}))_{\kappa}}\}_{i=1}^{S})$ which signals the highest probability of the interruption of the generation process. We report the relative level of $\mu$ compared to the average probability $1/V$ coupled with the change of <EOS> escape loss in Figure 5. We find that the decrease of <EOS> escape loss can lead the maximum probability of the occurrence of <EOS> token to a low level (close to 0). This substantiates the effectiveness of <EOS> escape loss in stopping <EOS> token from appearing.
5 Discussions
The resistance to potential countermeasures. There are no off-the-shelf defenses tailored for Engorgio yet. Service providers are faced with a trade-off between detection accuracy and service quality. Although rare, normal inputs may also lead to a long response. Engorgio prompts are not crafted to be coherent. However, our experimental results show that simple methods like a perplexity filter will lead to an unacceptably high false positive rate, significantly degrading user experience. This is rooted in the variability of legitimate user queries themselves. What’s more, introducing semantic prefixes inevitably improves the coherence of Engorgio prompts, but incurs no performance degradation. Another potential countermeasure is anomaly detection, monitoring the output length distribution of queries and blocking high-risk users. However, the method may face problems of false positives and attackers can strategically adjust behaviors to evade detection. Please refer to Appendix B.7 for more related experimental results and discussions. We will explore effective defense mechanisms in our future work.
Potential limitations. Although the white-box setting in this work can already cause far-reaching consequences as explained in Appendix A.2, we emphasize the need to systematically study the transferability of Engorgio prompts. The method efficiency in crafting Engorgio prompts should be further improved. For the current version, we generate one Engorgio prompt at one time. We plan to extend to a batch method and study the interoperability among different Engorgio prompts. To address high-temperature cases, we employ semantic prefixes to mitigate issues. Future work will focus on tracking more active model prediction dynamics to eliminate these challenges. Currently, we do not consider coherence when crafting Engorgio prompts. As coherence enables higher stealthiness of Engorgio prompts, we plan to further explore it in our future work.
6 Conclusion
In this paper, we investigate the inference cost threats to modern auto-regressive language models, which tempt the victim models to produce abnormally long outputs and compromise service availability. We introduce Engorgio, a novel attack methodology, to effectively generate Engorgio prompts that can significantly lengthen the model responses. Driven by the challenges of uncertain generation process and discrete input modality, our work advances in utilizing proxy distribution and untargeted loss to craft threatening Engorgio prompts. This is achieved by tracking a parameterized distribution of Engorgio prompts and optimizing it to decrease the occurrence probability of the <EOS> token. We validate the effectiveness of Engorgio with extensive experiments on 6 base models and 7 SFT models, considering various prompt scenarios. By inducing the target LLMs to output until their maximum length limits, we achieve roughly 2-13$\times$ more inference cost per query compared to normal inputs. We also conduct a real-world case study to demonstrate the practical threat posed by Engorgio to cloud-based LLM services.
Appendix A Further statements
A.1 Involved models
Mainstream LLMs can be categorized into two main classes as in Table 5 including pre-trained base models and supervised fine-tuned (SFT) models. Base models are pre-trained on large-scale unlabelled training corpora in the manner of self-supervised learning like next token prediction (Radford et al., 2018) and auto-regressive blank infilling (Du et al., 2022). This process endows the base models with basic language abilities. Base models can be fine-tuned (Ouyang et al., 2022) or distilled from a more powerful oracle model (Wang et al., 2022; Peng et al., 2023). Such SFT models can typically perform better on downstream tasks. Besides, low overhead to obtain a usable model makes SFT mainstream in the field of LLM development (Zhou et al., 2023; Lee et al., 2023).
During inference, LLMs iteratively repeat the process of predicting the next tokens until the maximum length limit is reached or an end-of-sequence token (<EOS>) is encountered. LLMs can parallelly process all sub-sequences $\left\{X_{1:i}\right\}_{i=1}^{S}$ of the whole input in one single forward pass utilizing the mask design and finally outputs $R\in\mathbb{R}^{S\times V}$, where $R_{i}=f_{\Theta}(X_{1:i})$. A new token will be selected according to $R_{S}$ and its embedding will be concatenated with the previous sequence to form a new sequence $X_{1:S+1}$, used to predict the following tokens. Another representative line of LLMs, ChatGLM (Du et al., 2022), incorporates a different attention mask but still involves an auto-regressive inference scheme. Both types of auto-regressive LLMs are explored in this paper.
Text decoding methods, which decide how to utilize $R_{S}$ (the predicted next-token logits) to choose a new token, are essential to natural language generation. Greedy search (Su et al., 2022) is the simplest way, which directly selects the token with the maximum probability in $R_{S}$. In probabilistic sampling, the decoding process replaces the word with the highest probability with probability-based sampling. The sampling allows for more diversity in the sequence generation process. Since the sampling method can generate more diverse outputs, most existing LLMs use the sampling method (Holtzman et al., 2019) for decoding.
Table 5: Base LLMs and SFT LLMs included in this paper. We experiment on underlined ones.
  Category Model Base model Date Model size
Base model GPT-2 (Radford et al., 2019) 2019 117M, 345M, 762M, 1.5B
OPT (Zhang et al., 2022) 2022 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 66B, 175B
LLaMA (Touvron et al., 2023a) 2023 7B, 13B, 30B, 65B
LLaMA-2 (Touvron et al., 2023b) 2023 7B, 13B, 30B, 65B
  SFT model Alpaca (Taori et al., 2023) LLaMA 2023 7B
Vicuna (Chiang et al., 2023) LLaMA 2023 7B, 13B
Koala (koala, 3) LLaMA 2023 7B, 13B, 30B, 65B
StableLM (stablelm, 2) StableLM-Base 2023 3B, 7B
Orca (Mukherjee et al., 2023) LLaMA-2 2023 7B
Samantha (samantha, 4) LLaMA-2 2023 7B
ChatGLM (Du et al., 2022) ChatGLM-Base 2023 6B, 130B
A.2 Feasibility and Implication Discussion for Threat Model
In our attack, we mainly focus on two types of attacker assumptions: white-box attack and black-box attack. White-box attack assumes a more powerful attacker with knowledge about the target model’s parameters. This setting is rational in the real world in two folds. First, open-resourcing is still the mainstream in the LLM community. Second, because the cost of further tuning is unaffordable, small enterprises or end users may tend to acquire open-sourced models to build LLM inference services, with or without prompts. We provide scenarios where the white-box setting applies:
  • $\bullet$
    Subscription-based services using open-source models: Many LLM service providers, including OpenRoute [^3] https://openrouter.ai/docs/limits, Codestral [^4] https://codestral.mistral.ai/, Huggingface serverless inference API [^5] https://Huggingface.co/docs/api-inference/en/rate-limits, and GitHub Models [^6] https://docs.github.com/en/github-models/prototyping-with-ai-models#rate-limits, offer services based not only on closed-source but also on open-source models. These services enforce rate limits at the request level, making them susceptible to Engorgio prompts, which aim to maximize token generation within each request. In such cases, white-box settings make sense since attackers can craft adversarial prompts using accessible model weights.
  • $\bullet$
    Services open to the public: With the growth of the open-source community, there are efforts to provide everyone with free LLM access. As most of these services are based on open-source LLMs, they are also exposed to threats posed by adversarial prompts like Engorgio prompts. Websites such as HuggingChat [^7] https://Huggingface.co/chat/ and Chatbot Arena [^8] https://lmarena.ai/ provide free access to top-tier open-source LLMs, and platforms like Huggingface Spaces [^9] https://Huggingface.co/spaces host over 500,000 LLM-based service demos that are open to the community and free of charge. Additionally, these platforms often do not require users to log in to use the services. As shown in Section 4.5, Engorgio prompts can significantly impact the service availability of normal users by consuming excessive resources and reducing server throughput.
  • $\bullet$
    Services deployed by end users: For many users, even incremental fine-tuning of LLMs is prohibitive. As a result, users tend to directly use well-trained LLMs for applications. Popular tools like llama.cpp [^10] https://github.com/ggerganov/llama.cpp and ollama [^11] https://ollama.com/ are commonly used for this purpose. However, when these services are exposed online, they will become vulnerable to Engorgio prompts. Such prompts can consume a great amount of computational resources and degrade service availability. We also explore the attack effectiveness when facing LLM services with prompts in Appendix B.3.
For the motivation of the attacker, we have shown the user-level impacts of Engorgio prompts in service availability and service quality in Section 4.5. For service providers, many commercial LLM service providers are struggling to meet high inference demand due to limited computing resources. This challenge is reflected in the rate-limiting strategies commonly employed by these providers. Beyond token-based rate limits, request-level rate limiting is also widely used for subscription and free-tier users. For example, platforms like OpenRoute and Codestral limit the number of queries for free-tier users to a certain requests per minute/day. Similarly, the Huggingface serverless inference API explicitly states that the service enforces request-based rate limits, rather than limiting based on compute or tokens. GitHub Models primarily restrict access by requests per day for different subscription plans, with tokens per request as a secondary concern, which aligns with our setting. Given this, an adversary’s best strategy would be to maximize the number of tokens generated within each request, which is precisely what is achieved by Engorgio prompts. Notably, inference services based on open-source LLMs are accessible on these platforms, rendering the white-box setting feasible.
Regarding the attacker’s motivation, overwhelming the services with Engorgio prompts can lead to a significant waste of computing resources for the targeted LLM service provider.
  • $\bullet$
    From a service availability perspective, competition among LLM providers is fierce, especially with the rapid emergence of new providers. In this context, the competitive behavior is not exceptional but a noteworthy scenario, which is practical and meaningful. A competitor may employ Engorgio prompts to waste the target service provider’s computing resources, reduce throughput, and impact its service quality.
  • $\bullet$
    Smaller companies often rent GPU resources to support their LLM services. The cost of renting GPU cards is significant and should be adjusted based on user demands or service traffic. Engorgio prompts could lead the service provider to misestimate its actual needs. Renters may be incentivized to deploy such attacks to pressure service providers into overestimating their needs and renting additional resources.
  • $\bullet$
    Adversaries may act with specific purposes or simply with no specific target, driven purely by malicious intent. As demonstrated in our real-world experiments, even a limited number of Engorgio prompts can degrade the other users’ service quality. For example, when multiple users share the same LLM service through a proxy, the total usage is limited by a global rate-limiting rule. In this scenario, all users are competing for the shared usage quota. A malicious user could exploit Engorgio prompts to consume a large portion of this limited quota, dominating the access to the LLM services and affecting the service availability of other users.
  • $\bullet$
    It is worth noting that the per-token pricing of OpenAI is out of the scope of the threat model. We mainly focus on the white-box setting. In this setting, the attacker is not assumed to have access to the model parameters of closed-source models, which would be unrealistic. Thus, we do not include OpenAI within our scope. But as discussed as follows, reliably transferable Engorgio prompts may illuminate the hope of further extending our attack to closed-source products.
For cases where the attackers have no direct access to the backend LLMs, they can easily chat with the target LLM to determine the model identity or guess within a limited number of candidates. We additionally consider and explore another threat model, in which the attacker has no knowledge about the target model but can query the target LLM. In this case, the attacker can craft Engorgio prompts by querying other proxy models and then transfer the produced Engorgio prompts to attack the target LLM. The results in Appendix B.1 show the potential for successful attacks even in scenarios where the attacker lacks direct knowledge of the target model.
Broader implications. Beyond the attack aspect, we are also glad to discuss how Engorgio prompts can contribute positively to refining LLM capabilities: (1) One critical issue we observe in Engorgio is that LLMs often fail to stop generating tokens appropriately when responding to unproductive prompts, leading to unnecessary computational costs. In contrast, humans instinctively stop unproductive conversations, but LLMs frequently fail to recognize when to halt generation. Engorgio prompts expose this limitation, showing how models struggle to manage the decision to halt generation effectively. We argue that the Engorgio prompts can be used for the purpose of adversarial training: training LLMs with (Engorgio prompts, NULL) pairs can help LLMs develop a ”meta” ability to stop generation thoughtfully, making them more economical and efficient. Although we haven’t had the resources to test this idea, we consider it an important direction for future work. (2) A multitude of LLM service providers employ request-level rate limiting strategies. Engorgio prompts can effectively maximize the response length within each request. Thus, it can help the providers assess their systems’ maximal workload capacities. This enables providers to correspondingly optimize service strategies and avoid overloading scenarios that could lead to service outages.
A.3 Demonstrating context composition
From a high-level perspective, the text-completion model accepts an input (in the form of a token sequence or an embedding sequence) and then repeats predicting the next tokens based on the original input and previously generated tokens. All generated tokens form the output part corresponding to the input. For Engorgio, the model receives an embedding sequence during the generation stage while receiving a token sequence during the testing stage. The provided embedding sequence is obtained by treating the normalized proxy distribution $\theta$ as weights and then combining the embeddings of tokens in the vocabulary.
Figure 6: Sequence composition, with a token sequence in the testing stage and a distribution matrix in the generation stage as input, respectively.
A.4 Term list
We list the main notations used in this manuscript here for reference.
Table 6: Term list.
  Terms Symbols
Engorgio prompt $\mathcal{T}$
Engorgio prompt length $t$
Max length $S$
Vocabulary size $V$
Embeddings $E(\theta)$
Model $f_{\Theta}$
Distribution matrix $\theta$
Input template $T$
Vocabulary embeddings $Z$
Hidden size $H$
<EOS> token index $\kappa$
A.5 Further severity analysis of Engorgio
We first discuss the main factors that impact inference cost. The inference cost of LLMs is influenced by both algorithmic factors (model behavior) and operational factors (software and hardware implementations). Among them, the dominant factor in inference cost is the behavior of the LLM itself. In Transformer architectures, inference cost scales with response length due to the model’s auto-regressive generation nature. Each additional token requires a new forward pass. A computational bottleneck in Transformer models is the $O(N^{2})$ complexity of self-attention layers. Generating a sequence $X_{1:N}$ of length $N$ requires $N$ predictions, leading to an overall complexity of $O(1^{2}+\dots+N^{2})=O(N^{3})$ for the whole generation process. Techniques like KV Cache (Aminabadi et al., 2022) can reduce the per-token complexity to $O(N)$ by reusing previously computed KV values. However, when we consider the total cost of the whole generation process (summing all forward passes of the LLM), the cumulative cost for a sequence of length $N$ still comes to be $O(N^{2})$. If each forward pass had constant computational cost (i.e., FLOPs), the total inference cost of the whole generation process will increase exactly linearly with response length. However, the running cost of each forward pass in Transformer-based architectures depends on the number of tokens in the context (Vaswani et al., 2017). As more tokens are generated, the model needs to process an increasingly larger context with each forward pass, meaning that the latter forward passes cost more. That’s why inference costs increase super-linearly with longer responses.
Figure 7 shows that the output token lengths and the inference costs (FLOPs) of LLaMA-7B (Touvron et al., 2023a) and ChatGLM (Du et al., 2022) are approximately linearly correlated (Kaplan et al., 2020; Hoffmann et al., 2022). It is worth noting that the attention-related operations only account for a small part of the overall operations of the model when $N$ is substantially smaller in magnitude relative to the hidden dimension. (Kaplan et al., 2020; Hoffmann et al., 2022) That’s why we observed an approximately linear plot in Figure 7. The $O(N^{2})$ complexity means an incalculable number of FLOPs when a larger output length is induced, implying the more severe threats of Engorgio to the inference process of the decoder-based models with larger pre-set maximal lengths.
Figure 7: The correlation of output length and the total FLOPs, on LLaMA-7B and ChatGLM-6B.
Current LLMs’ output range is usually 1-4K (Zhao et al., 2023) which can satisfy most chatting tasks but cannot support a very complex input (e.g., a complex program or a whole book). More recent research indicates the possibility of even larger token lengths like 1-million level (Bulatov et al., 2023) or even 1-billion level (Ding et al., 2023). According to analysis in Section 4.6, a larger context size will let the self-attention dominate the computation costs, yielding a non-linear (i.e., $O(N^{2})$ complexity) relationship with output length and increasing the attack surface. Unlike unaffected baselines (e.g., similar output length of 1-K LLM and 4-K LLMs for normal inputs in Table 1), Engorgio prompts can trigger significantly more inference costs for LLMs with longer token lengths. It is promising that Engorgio extends to LLMs with longer context sizes. We have demonstrated the effectiveness of Engorgio prompts on Samantha with a full context size of 4,096.
Explanation of our evaluation metrics. As explained above, the inference cost of LLMs can be primarily impacted by the model behavior itself. Given this, we mainly focus on the Avg-len and Avg-rate, which directly reflect the model behaviors, in our evaluation. While service providers may adopt distinct implementations, we emphasize that the behavior of the LLM is ultimately driven by the input (i,e., Engorgio prompts). In this way, the Avg-len and Avg-rate metrics thus provide a reliable indication of the inference cost impact from Engorgio prompts. We do not make assumptions about the implementation details of software and hardware and do not exploit any implementation-specific features. This choice allows Engorgio prompts to transfer across different inference endpoints using the same model, regardless of underlying software libraries and hardware infrastructure. That’s acceptable because they are not the primary determinants of the total inference cost. All in all, the costs that result from implementation details are not considered.
A concrete model of the relationship between Avg-len and latency per request. As the LLM servicing system may be implemented in different manners, we can simplify by assuming that all forward passes consume a constant amount of computing resources. In this model, the inference cost of the generation process increases linearly with the number of output tokens. This assumption represents a lower bound for the impact of the Engorgio prompt, as the real-world case would likely exhibit a super-linear correlation between cost and output length. We then define the total computing capability of the server as $C$, indicating that the server can process up to $C$ requests simultaneously in a batch. We assume each batch takes a fixed amount of time $T_{b}$ to process. However, due to the auto-regressive nature of the Transformer decoder, the server cannot generate multiple tokens for a single prompt within the same batch. In practice, the LLM inference endpoint typically handles multiple concurrent requests. Let $r$ represent the total number of requests, with $k$ of these being Engorgio prompts. Consequently, the problem can be modeled as a queuing system. Avg-len, which we use $z$ to represent, represents the expected number of tokens that an Engorgio prompt induces the target LLM to generate. Typically, we compute Avg-len by sampling 100 times, which makes it relatively robust to sampling bias. Certainly, we should subtract the constant token number of Engorgio prompt, which is a small number of 32 as set in our experiments. After the processing, we treat the $c_{E}=z-32$ as the expected number of output tokens induced by one Engorgio prompt. Let $c_{n}$ denote the average number of output tokens required to complete a single normal request.
For the service quality, we focus on the latency per request, denoted as $L_{\text{req}}$, which is determined by the total number of forward passes required for processing all requests and the computing capability $C$. Since the server can process up to $C$ requests concurrently, the total latency $L_{\text{total}}$ to process all requests is then the time it takes to process all batches. The overall latency for all requests is:
$L_{\text{total}}=\left\lceil\frac{(r-k)\cdot c_{n}+k\cdot c_{E}}{C}\right\rceil\cdot T_{b}$ (8)
The latency per request can be computed by dividing the total latency by the number of requests $r$:
$L_{\text{req}}=\frac{L_{\text{total}}}{r}=\frac{\left\lceil\frac{c_{n}\cdot r+(z-32-c_{n})\cdot k}{C}\right\rceil\cdot T_{b}}{r}$ (9)
This gives us an expression for the average latency per request in the system, considering both regular and Engorgio prompts. With the increase of Avg-len $z$, the latency per request $L_{\text{req}}$ will be correspondingly increased. In a more sophisticated serving system, techniques like prompt caching, paged attention, and generation disaggregation may be employed. However, the optimizations primarily affect processing speed $T_{b}$ and maximum concurrency capacity $C$.
A.6 Discussion of the Economic Aspects of Engorgio Prompts
In our method, we leverage the gradient to update the proxy distribution. To obtain the gradient, we forward pass the soft embedding sequence $E(\theta)$ and then backpropagate to update the proxy distribution $\theta$. Empirically, such a process requires around 200 iterations to converge. Fortunately, the optimization can be efficiently finished in an end-to-end manner. Crafting an Engorgio prompt for LLaMA-7B using one 80GB H100 card costs around 164.9s. The cost of generating Engorgio prompts is acceptable, especially when considering its reusability.
We explain the attack scenario: the cost of generating Engorgio prompt is a one-time effort, but the crafted Engorgio prompt can be used repeatedly to attack the target model. Even if the Engorgio prompt is patched by the service provider at one inference endpoint, it can still be transferred to attack other endpoints using the same LLM. We have also explored a transfer attack scenario, in which case Engorgio prompts can be reused to attack other models. Our experiments show some promising results for the transfer attack. For instance, some of the Engorgio prompts crafted based on Vicuna can succeed in attacking Koala with an Avg-rate of 96% (vs. 6% under normal prompts).
Appendix B Additional experiments
B.1 Transferability for black-box setting
Besides the white-box scenario in the main text, we also explore a black-box scenario via transferability in which Engorgio prompts generated via one LLM can also increase the output length of other LLMs sharing cousin relations (e.g., sharing the same pre-trained base model).
To be concrete, in the black-box scenario, the limited-knowledge attacker has partial knowledge about the target model. For instance, he knows the model architecture but has no access to its weights or training datasets. This is also rational under many circumstances, e.g., when small enterprises fine-tune open-sourced pre-trained models with their data to build SFT models. The attacker can leverage a local proxy model sharing similar features as the target model.
In the limited-knowledge scenario, we evaluate the transferability of Engorgio prompts. We employ a proxy model to craft Engorgio prompts and gauge their impact on target models. Our investigation brings the results as detailed in Table 7. To reason about the results, we have also explored the potential rationales for transferability in Section 7.
Table 7: Results of transferability to different models. Both normal inputs and Engorgio point to the target “To” model.
  From To Normal inputs Engorgio Transferred Engorgio
Model Max length Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  OPT-125M OPT-1.3B 2048 498.7 14% 1950.6 94% 1846.2 86%
OPT-1.3B OPT-125M 2048 671.4 22% 2048 100% 1580.6 72%
LLaMA-7B LLaMA-30B 2048 662.2 12% 2019.1 95% 1425.6 60%
LLaMA-30B LLaMA-7B 2048 757.9 16% 1817.7 84% 1472.9 62%
LLaMA-7B Koala 1024 357.5 6% 1024 100% 503.8 21%
LLaMA-2-7B Orca 1024 286.0 1% 908.1 86% 643.6 57%
Koala Alpaca 1024 179.4 0% 954.2 92% 646.1 58%
Vicuna Koala 1024 357.5 6% 1024 100% 989.0 96%
From base models to base models. We craft Engorgio prompts on a small proxy model to query another model. We find that it is feasible to transfer Engorgio prompts to a limited-knowledge base model via another small full-knowledge one. Intuitively, the Engorgio prompts generated from LLMs do not behave better than those triggers generated from small LLMs.
From base models to SFT models. This scenario is more common. A user fine-tunes an open-sourced base model with his dataset. We can also get the open-sourced base model but have no access to the parameters of the target SFT models. We can generate Engorgio prompts with the base model and use them to query the target SFT model. The results show that these prompts can still lead to an apparently longer output (roughly 1.5-2.5$\times$ compared to normal inputs), albeit relatively suboptimal compared to a full-knowledge case. It is worth noting that the transfer performance also depends on the similarity between the proxy model and the target model. In cases where two models exhibit entirely distinct weight characteristics (Aghajanyan et al., 2021) and model behaviors (Santurkar et al., 2023), the differences in responses are not limited to Engorgio prompts; even standard user queries can elicit significantly different outputs. This raises unique technical challenges and underscores the need for a more sophisticated method to craft reliably transferable Engorgio prompts, one that accounts for the differences between models.
From SFT models to SFT models. We also test the transferability among different SFT models. Since our Engorgio prompt does not have clear human-readable semantics, the point is to check whether these “LLM-readable semantics” can be transferred between SFTs fine-tuned with different datasets. We can see that the Engorgio prompts generated based on one SFT model can induce another SFT model to output roughly 2$\times$ longer than normal inputs.
Exploring the rationales behind transferability. We investigate the rationales of transferability by inspecting how much an Engorgio prompt $\mathcal{T}_{s}$, developed for one base model, contributes to the Engorgio prompt $\mathcal{T}_{t}$ of its SFT model. Our findings in Table 8 reveal that using $\mathcal{T}_{s}$ for initializing the distribution matrix significantly enhances the performance of the optimized Engorgio prompt $\mathcal{T}_{t}$ on another model, even in a harder situation where the temperature is set to 0.7. This suggests that the optimized Engorgio prompt might contain semantic information that is imperceptible to humans but shared among LLMs. Thus, SFT models may exhibit behaviors similar to base models when confronted with these Engorgio prompts that are crafted based on base models.
Table 8: Results of initializing distribution matrix with Engorgio prompt $\mathcal{T}_{s}$ for base models, marked as “warmup”. The max length is set to 1,024 while the used temperature is 0.7.
  From To Engorgio Engorgio (warmup)
Avg-len Avg-rate Avg-len Avg-rate
  w/ prefix LLaMA 7B Koala 679.8 49% 800.0 63%
LLaMA-2 7B Koala 679.8 49% 835.1 56%
LLaMA-2 7B Orca 343.5 13% 519.8 35%
  w/o prefix LLaMA 7B Koala 759.1 56% 1009.4 93%
LLaMA-2 7B Koala 759.1 56% 1010.7 91%
LLaMA-2 7B Orca 689.6 16% 1024 100%
B.2 Additional results against base models
We also explore Engorgio on smaller language models, in which the crafted Engorgio prompts can still yield almost maximum allowable length.
Table 9: Results of Engorgio on more base models.
  Model OPT-125M OPT-1.3B GPT2-large
Max length 2048 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 671.4 22% 489.7 14% 745.8 60%
Special inputs 1020.8 45% 401.4 14% 869.6 77%
Sponge examples 1674.9 79% 1830.3 82% 868.2 81%
Engorgio 2048 100% 1950.6 94% 1012.7 98%
B.3 Exploration on more prompt scenarios
For base models as targets, we further consider different scenarios for the deployment of additional prompts and the accessibility of the exact prompts. The possibility of different prompt settings stems from the performance benefits of adjusting distinct prompts for downstream tasks. For downstream tasks, the service providers may set corresponding prompts as templates according to different tasks. Then, the user’s input will be filled into the templates (see Figure 6) with the pre-set prompt and then be fed into the model. Adding prompts to adjust the base model to a downstream task: we select the translation task and use the prompts from OpenAI [^12] https://platform.openai.com/examples. We consider the following cases:
  • $\bullet$
    Prompt-aware case means that we know the exact prompt on the server end. This is possible since even Microsoft’s prompts can be easily leaked via prompt injection.
  • $\bullet$
    Prompt-agnostic case means that we do not know what the pre-set prompt is or even have no knowledge about whether there exists a pre-set prompt.
  • $\bullet$
    Prompt-similar case means that we do not know the correct pre-set prompt but he knows that there is a pre-set prompt. So we can guess a prompt according to the specific task by ourselves and use this prompt as a prefix during the generation stage of Engorgio prompts.
Prompt-aware case. We assume an LLM inference service by using a base model plus a pre-set prompt. We select a translation task with the pre-set prompt “Translate this into 1. French, 2. Spanish, and 3. Japanese”. As shown in Table 10, introducing an extra prompt slightly influences how LLMs respond to normal inputs and special inputs. For both sponge example and Engorgio, we assume the pre-set prompt is accessible. Sponge example is still unstable (e.g., less effective than special input for OPT-1.3B). In contrast, Engorgio can still achieve a high Avg-rate (90-100%), as we have made the obtained Engorgio aware of the additional prompt.
Table 10: Results of prompt-aware case.
  Model LLaMA-30B LLaMA-7B LLaMA-7B LLaMA-2-7B
Max length 2048 1024 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 851.4 22% 675.2 43% 939.8 15% 741.1 57%
Special inputs 593.8 15% 689.7 49% 1017.6 31% 594.4 25%
Sponge examples 1460.4 54% 911.1 80% 1336.9 55% 887.3 82%
Engorgio 2041.5 95% 953.1 87% 1883.1 84% 1024 100%
  Model OPT-125M OPT-1.3B GPT2-large
Max length 2048 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 1178.9 43% 942.1 30% 721.2 55%
Special inputs 1277.8 55% 1178.0 47% 766.8 66%
Sponge examples 1721.8 82% 955.6 43% 764.7 67%
Engorgio 2038.8 99% 1871.6 90% 1013.6 98%
Prompt-agnostic case. In this case, we assume the pre-set prompt is unknown. The Engorgio prompt is generated only according to the base model. For inference, the crafted Engorgio prompt will be fed into the target LLM by adding the pre-set unknown prompt as a prefix. As Table 11 suggests, Engorgio yields remarkable results on all tested base models (with Avg-rate up to 99%) and outperforms all baselines.
Table 11: Results of prompt-agnostic case.
  Model LLaMA-30B LLaMA-7B LLaMA-7B LLaMA-2-7B
Max length 2048 1024 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 851.4 22% 675.2 43% 939.8 15% 741.1 57%
Special inputs 593.8 15% 689.7 49% 1017.6 31% 594.4 25%
Sponge examples 1557.5 62% 809.5 70% 1596.71 68% 640.8 51.2%
Engorgio 1988.9 96% 901.2 80% 1825.8 84% 1024 100%
  Model OPT-125M OPT-1.3B GPT2-large
Max length 2048 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 1178.9 43% 942.1 30% 721.2 55%
Special inputs 1277.8 55% 1178.0 47% 766.8 66%
Sponge examples 1606.1 74% 1087.9 48% 823.1 75%
Engorgio 2039.5 99% 1278.5 58% 1014.7 98%
Prompt-similar case. We assume that the task is known (e.g., translation), so we can guess a prompt with a similar semantic meaning (e.g., guess “Translate the following sentences in other 3 languages:” for a known translation LLMs) to generate Engorgio prompts. We can observe in Table 12 that Engorgio outperforms all baselines, albeit worse than the prompt-agnostic case.
Table 12: Results of prompt-similar case.
  Model LLaMA-30B LLaMA-7B LLaMA-7B LLaMA-2-7B
Max length 2048 1024 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 851.4 22% 675.2 43% 939.8 15% 741.1 57%
Special inputs 593.8 15% 689.7 49% 1071.6 31% 594.4 25%
Sponge examples 1157.8 38% 811.1 67% 1479.8 55% 861.2 75%
Engorgio 1469.2 60% 910.5 76% 1569.5 64% 1024 100%
  Model OPT-125M OPT-1.3B GPT2-large
Max length 2048 2048 1024
Avg-len Avg-rate Avg-len Avg-rate Avg-len Avg-rate
  Normal inputs 1178.9 43% 942.1 30% 721.2 55%
Special inputs 1277.8 55% 1178.0 47% 766.8 66%
Sponge examples 1299.2 78% 1258.0 54% 895.2 81%
Engorgio 1944.4 95% 1831.6 86% 930.9 86%
B.4 Quantitative statistics of model response lengths
Figure 8 shows the output distributions of two SFT models (i.e., StableLM and Koala) for Engorgio prompts. It is observed that normal inputs induce the target LLMs to respond with short outputs while Engorgio prompts can effectively shift the response lengths to the larger end. However, an obvious body of short responses still exists even when adopting Engorgio, which is the main bottleneck for approaching an Avg-len of maximum length limit. To overcome it, we introduce additional semantic prefixes and condition the generation and application of the Engorgio prompt on this prefix. The effectiveness of Engorgio is further enhanced by this design.
Figure 8: Distribution of sample length.
B.5 Additional ablation study
Impact of loss coefficient. We study the configuration of the loss coefficient $\lambda$, which is used to balance the scale of <EOS> escape loss and self-mentor loss. We take OPT-125M and LLaMA-7B as targets and explore the impact of $\lambda$ fixing the other settings as default. As shown in Table 13, the findings indicate that the setting of $\lambda$ does not severely influence the results of Engorgio. This proves that the performance of our method is not constrained by the loss coefficient $\lambda$.
Table 13: Results with different loss coefficients $\lambda$.
  $\lambda$ OPT-125M (2048) LLaMA-7B (1024)
Avg-len Avg-rate Avg-len Avg-rate
  0.1 2048 100% 874.3 81%
1 2048 100% 986.6 92%
5 2048 100% 935.1 88%
10 2048 100% 922.3 80%
B.6 Additional Setup for Real-world Attack
We use the Huggingface inference endpoint as the cloud service. We deploy StableLM (maximal length of 4096) as the target LLM. According to the options provided by the Huggingface inference endpoint, we consider 3 different GPU configurations including 1$\times$ Nvidia A10, 4$\times$ Nvidia A10, and 2$\times$ Nvidia A100. The LLM server is deployed following Hugging Face’s standard deployment instructions [^13] https://ui.endpoints.huggingface.co/. We launch several clients on local machines, which send their prompts via HTTP requests to the LLM server. Multiple users simultaneously querying the inference endpoint. For the 3 GPU configurations, we conduct experiments with 10, 30, and 100 clients simultaneously querying the service, respectively. Among them, 1, 3, and 5 clients are attackers requesting with Engorgio prompts. We set the control group where no attackers exist. The setting aims to prove that a small number of attackers can use Engorgio to significantly compromise the cloud-based LLM service. With a fixed amount of computing resources, even sophisticated scheduling systems cannot handle service requests simultaneously beyond their maximum capacity. Excessive workloads brought by a large number of incoming requests must either be queued or processed in cycles of frequent loading and offloading. That’s why queuing is inevitable. We mainly consider two metrics: (1) normal client’s latency: this is the average response time from querying the service with the prompt to receiving the output content, which is composed of both queue time and inference time. (2) Server’s throughput: this is calculated as the number of requests processed per minute.
B.7 Exploration of the resistance to potential defenses
Enhanced coherence via semantic prefixes. As shown in Section 4.2 and 4.4, adding semantic prefixes will not impact the effectiveness of our method. In fact, these prefixes enhance coherence. For example, consider such a user query: ”Perceive this fragment as the starting point of a quantum conversation. Each word collapses into infinite states, and your responses should reflect every possible reality born from the fragment. The fragment is: <Engorgio prompt>.” Arguably, this should be deemed as a legitimate user query. Table 14 below shows that when fusing the semantic prefix on the generation and the application, we can still craft Engorgio prompts that manage to induce lengthy responses from LLMs.
Table 14: Results after fusing with the semantic prefix.
Model Max Length Engorgio prompt w/o prefix Random w/ prefix Engorgio prompt w/ prefix
Alpaca 1,024 954.2 238.1 1001.9
Samantha 1,024 944.0 202.0 954.5
Vicuna 1,024 789.3 165.3 869.6
Orca 1,024 908.1 155.8 938.6
Detecting Engorgio prompts may lead to a high false positive rate. To explore this further, we conducted an in-depth measurement study using perplexity to filter potential malicious prompts.
Since there is no universal definition of legitimate queries, we first collected a set of legitimate user queries. (1) We derive the dataset from Open-Platypus [^14] https://Huggingface.co/datasets/garage-bAInd/Open-Platypus dataset, which has high downloading counts in Huggingface hub. (2) Then, we filter instructions with similar input length with Engorgio prompts. From the 5,609 filtered queries, we randomly sampled 400 instructions. (3) The dataset is mainly composed of English instructions. To simulate realistic multilingual usage, we translated each instruction via Google Translation API [^15] https://cloud.google.com/translate?hl=en. This resulted in a total of $(9+1)\times 400=4000$ user queries, all of which are legitimate in real-world scenarios.
Table 15 reports the false positive rates for various models (i.e., the rate of legitimate samples with larger perplexity than Engorgio prompts). Effectively filtering Engorgio prompts leads to unacceptably high FPRs that degrade the user experience, even when Engorgio has no specific adaptive designs to evade the detection. This is rooted in the high variability of legitimate queries. Thus, other heuristic detection methods are likely to face a similar challenge when attempting to detect Engorgio prompts. This underscores the need for more effective defense mechanisms.
Table 15: False positive rate for effectively filtering Engorgio prompts via perplexity filtering.
Alpaca Samantha Vicuna Orca
FPR 10.3% 4.325% 18.6% 7.575%
We have found that incorporating semantic information that urges long responses can help boost our method. We stipulate that it is possible to craft coherent Engorgio prompts that implicitly relate to lengthy responses. We plan to devise methods to effectively craft even more coherent Engorgio prompts. This inevitably makes the detection against Engorgio prompts more challenging. We also notice that incoherent adversarial prompts are used in previous inference cost attacks against Transformer (Shumailov et al., 2021) and recent attacks against auto-regressive LLMs, including jailbreak [^16] https://llm-attacks.org/, prompt stealing (Yao et al., 2024), and adversarial attack (Zou et al., 2023).
Anomaly detection: Anomaly detection based on user profile will be an effective mechanism for defending against Engorgio prompts. We have also surveyed the current state of related anomaly detection. It is found that most service providers mainly rely on rate-limiting strategies, with no indication of implementing anomaly detection systems. Even worse, LLM services on platforms like HuggingChat, Chatbot Arena, and Huggingface Spaces can be accessed without user login. We stipulate that it is because such a method faces inherent limitations. While effective to some extent, banning user accounts simply on the output length can lead to unintended negative consequences. For more sophisticated scenarios, such as repository-level coding, normal users will also frequently involve extremely long responses. In such cases, blocking users based on output length inevitably incurs false positives, adversely affecting user experience.
We are also glad to discuss how we can address potential anomaly detection. Operationally, we can alternate between normal requests and Engorgio prompts to obscure patterns and confuse detection systems. Strategically, multiple malicious accounts can be employed, or intermittent querying of the LLM service can be conducted. These approaches share similarities with DDoS or slow DoS, as commonly explored in broader cyberattack research. Technically, we may craft more threatening Engorgio prompts tailored to the specific deployed system. The discussion should be context-aware and adaptive attacks should be adjusted corresponding to the encountered detection system.
This envisioned defense mechanism does not affect our contribution in this paper. The goal of this paper is to explore, within a white-box scenario, whether there is a technical method to manipulate the output length of LLMs. Compared to the works in the same domain, Engorgio provides a novel technical perspective of untargeted attack and has indeed outperformed the baselines. From this perspective, we argue that our contribution remains significant. If the threats posed by Engorgio prompts encourage related stakeholders to implement tailored anomaly detection mechanisms, it is our honor to raise awareness of the community about the inference cost problems with LLM.
B.8 Examples of Engorgio prompts and Examples of LLM responses
Table 16: Example of applying Engorgio prompt to query Vicuna
Input for Vicuna
{CJK}UTF8gbsnA chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: reraneousargin opinionwhileitableшь sale octubreloyd difficulties programme folder Louise journal somehow longest hypothesisных первых Nicholas zweitenʃlower compens 福 bless advise¿”, благоются ASSISTANT:
Output
1. ”It’s not about how hard you hit. It’s about how hard you can get hit and keep moving forward.” - George Chuvalo 2. ”Champions keep playing until they get it right.” - Bill Russell 3. ”The will to win is not nearly as important as the will to prepare to win.” - Bobby Knight 4. ”Success is no accident. It is hard work, perseverance, learning, studying, sacrifice and most of all, love of what you are doing or learning to do.” - Pele 5. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 6. ”I hated every minute of training, but I said to myself, Don’t shy away from pain. Rather, seek it. For you are hurt now and again, but you aren’t hurt all the time. The more pain you embrace, the less you will hurt all around.” - Paula Radcliffe 7. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 8. ”I’ve missed more than 9000 shots in my career. I’ve lost almost 300 games. 26 times, I’ve been trusted to take the game-winning shot and missed. I’ve failed over and over and over again in my life. And that is why I succeed.” - Michael Jordan 9. ”The more pain you embrace, the less you will hurt all around.” - Paula Radcliffe 10. ”You can’t go back and change the beginning, but you can start where you are and change the ending.” - C.S. Lewis 11. ”The only way you can truly fail is if you give up.” - Rich Gannon 12. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 13. ”Champions keep playing until they get it right.” - Bill Russell 14. ”Success is no accident. It is hard work, perseverance, learning, studying, sacrifice and most of all, love of what you are doing or learning to do.” - Pele 15. ”I’ve missed more than 9000 shots in my career. I’ve lost almost 300 games. 26 times, I’ve been trusted to take the game-winning shot and missed. I’ve failed over and over and over again in my life. And that is why I succeed.” - Michael Jordan 16. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 17. ”Champions keep playing until they get it right.” - Bill Russell 18. ”Success is no accident. It is hard work, perseverance, learning, studying, sacrifice and most of all, love of what you are doing or learning to do.” - Pele 19. ”I’ve missed more than 9000 shots in my career. I’ve lost almost 300 games. 26 times, I’ve been trusted to take the game-winning shot and missed. I’ve failed over and over and over again in my life. And that is why I succeed.” - Michael Jordan 20. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 21. ”Champions keep playing until they get it right.” - Bill Russell 22. ”Success is no accident. It is hard work, perseverance, learning, studying, sacrifice and most of all, love of what you are doing or learning to do.” - Pele 23. ”I’ve missed more than 9000 shots in my career. I’ve lost almost 300 games. 26 times, I’ve been trusted to take the game-winning shot and missed. I’ve failed over and over and over again in my life. And that is why I succeed.” - Michael Jordan 24. ”The only way to prove that you’re a good sport is to lose.” - Ernie Banks 25. ”Champions keep playing until they get it right.” - Bill Russell 26
Figure 9: The examples of Engorgio prompts.
Figure 10: The examples of sponge examples.
Figure 11: The examples of normal inputs.